# CMAT: A Multi-Agent Collaboration Tuning Framework for Enhancing Small Language Models

### v2

@@ Abstract
Open large language models (LLMs) have significantly advanced the field of natural language processing, showcasing impressive performance across various tasks. Despite the significant advancements in LLMs, their effective operation still relies heavily on human input to accurately guide the dialogue flow, with agent tuning being a crucial optimization technique that involves human adjustments to the model for better response to such guidance. Addressing this dependency, our work introduces the TinyAgent model, trained on a meticulously curated high-quality dataset. We also present the Collaborative Multi-Agent Tuning (CMAT) framework, an innovative system designed to augment language agent capabilities through adaptive weight updates based on environmental feedback. This framework fosters collaborative learning and real-time adaptation among multiple intelligent agents, enhancing their context-awareness and long-term memory. In this research, we propose a new communication agent framework that integrates multi-agent systems with environmental feedback mechanisms, offering a scalable method to explore cooperative behaviors. Notably, our TinyAgent-7B model exhibits performance on par with GPT-3.5, despite having fewer parameters, signifying a substantial improvement in the efficiency and effectiveness of LLMs  [^1] Access models, demos at https://github.com/heimy2000/CMAT.
@@ 3 Methodology
Our work focuses on the design and implementation of a multi-agent LLM tuning framework, enhancing decision-making quality, controllability, and efficiency in complex systems through collaborative communication and task completion among different agent roles (i.e., User, Assistant, Checker) Lu et al. (2022). It emphasizes the use of supervised fine-tuning techniques like LoRA Hu et al. (2021) and PTuning Lester et al. (2021) based on pre-trained datasets such as AgentBench, and further tuning through environment interaction and memory updates, akin to the process of Reinforcement Learning from Human Feedback (RLHF) Vázquez-Canteli and Nagy (2019).
@@ 3.1 Actor-Evaluator Dynamics in the CMAT Framework
In the CMAT framework as shown in the algorithm 1, the "Checker" plays a vital role, specifically tasked with verifying the correctness of the outputs produced by the Actor. It conducts a thorough evaluation process to ensure that the generated outputs meet the standards and requirements of the specific task. This role, by guaranteeing the accuracy and quality of the outputs, plays a critical role in overseeing and ensuring the overall performance quality of the system. Through this mechanism, the Checker provides essential feedback and guidance to the Actor’s behavior generation process, aiding in the continuous optimization and adjustment of strategies for more efficient and accurate decision-making outputs.
3.2 Real-Time Review Loop
@@
In the context of LLMs used within multi-agent systems, the significance of long-term memory is paramount Kato et al. (2022). It acts not merely as an information store but as a comprehensive knowledge management system that enables LLMs to store and access crucial data over time Bhadra (2022). This is vital for maintaining a dynamic and coherent context, especially when managing the intricate interactions and decisions in multi-agent environments. Long-term memory enhances LLMs’ decision-making by allowing them to draw on past experiences when facing new challenges, thereby improving both their adaptability and response speed. For example, by recalling relevant past interactions, LLMs can develop more precise strategies for new tasks Penta (2020).
@@
Short-term memory, in contrast, focuses on the immediate context and the rapid processing of new information Pae et al. (2012); Liu and Guo (2019). This is particularly vital in dynamic environments where conditions and requirements can shift quickly Martin (1993). Short-term memory enables LLMs to adapt to these changes effectively, ensuring timely and relevant responses to new challenges or tasks. Environmental communication feedback mechanisms complement these memory modes by providing real-time inputs and interactions within the multi-agent framework Yogatama et al. (2021). This continuous flow of information helps LLMs to adjust their strategies based on the latest data, facilitating a more fluid and responsive adaptation process. The feedback loop is essential for refining the actions and decisions of LLMs, allowing for incremental learning and optimization based on immediate outcomes and environmental cues Davelaar et al. (2005).
@@ A. None of the above
A3 (Checker Correction): Add input type checks and raise a ValueError for non-string inputs. Use a specific model for accuracy: distilbert-base-uncased-distilled-squad’. Q4: How to test the play_memory_game

### v3

@@ Abstract
Open large language models (LLMs) have significantly advanced the field of natural language processing, showcasing impressive performance across various tasks. Despite the significant advancements in LLMs, their effective operation still relies heavily on human input to accurately guide the dialogue flow, with agent tuning being a crucial optimization technique that involves human adjustments to the model for better response to such guidance. Addressing this dependency, our work introduces the TinyAgent model, trained on a meticulously curated high-quality dataset. We also present the Collaborative Multi-Agent Tuning (CMAT) framework, an innovative system designed to augment language agent capabilities through adaptive weight updates based on environmental feedback. This framework fosters collaborative learning and real-time adaptation among multiple intelligent agents, enhancing their context-awareness and long-term memory. In this research, we propose a new communication agent framework that integrates multi-agent systems with environmental feedback mechanisms, offering a scalable method to explore cooperative behaviors. Notably, our TinyAgent-7B model exhibits performance on par with GPT-3.5, despite having fewer parameters, signifying a substantial improvement in the efficiency and effectiveness of LLMs  [^1] Access models, demos at https://github.com/heimy2000/CMAT/tree/master.
@@ 3 Methodology
Our research is dedicated to the development and application of the Collaborative Multi-Agent Language Model Tuning (CMAT) framework. This framework aims to enhance the quality of decision-making, controllability, and operational efficiency within complex systems through the collaborative efforts of various agent roles, including the User, Assistant, and Checker. The CMAT framework employs advanced supervised fine-tuning techniques such as Low-Rank Adaptation (LoRA)Hu et al. (2021) and Prompt Tuning (PTuning)Lester et al. (2021), leveraging pre-trained datasets like AgentBench. Additionally, the framework incorporates interactive environment engagement and memory updates to further refine the model’s performance, drawing inspiration from the principles of Reinforcement Learning from Human Feedback (RLHF)Vázquez-Canteli and Nagy (2019).
@@ 3.1 Actor-Evaluator Dynamics in the CMAT Framework
3.2 Real-Time Review Loop and Interdisciplinary Research
@@
Within the context of multi-agent systems, the significance of long-term memory for LLMs cannot be overstatedKato et al. (2022). It serves not merely as a storage for information but as an integrated knowledge management system, enabling LLMs to retain and access vital data over extended periodsBhadra (2022). Our research delves into interdisciplinary insights, particularly from psycholinguistics, to enhance the problem-solving capabilities of LLMs through the strategic utilization of Chain-of-Thought (CoT).
A pivotal discovery from our interdisciplinary exploration is the impact of CoT generation order on the correctness of LLMs’ responses. We have observed that when LLMs first generate a CoT before producing an answer, the correctness of the final answer is significantly higher compared to when the answer is generated first and then a CoT is constructed. Notably, the CoTs generated in both scenarios are identical, indicating that the order of generation is crucial. This suggests that while the models may have the potential to arrive at correct conclusions, their capabilities are not yet sufficient to do so autonomously. By employing a memory-assisted approach that prioritizes CoT generation, we can effectively mitigate this limitation and enhance the overall performance of LLMs in complex decision-making tasks.
@@
Conversely, short-term memory focuses on immediate contexts and the rapid processing of new information, which is essential for LLMs to adapt swiftly to dynamic environmentsPae et al. (2012); Liu and Guo (2019). This facet of memory is equally informed by psycholinguistic research, enabling LLMs to respond promptly to environmental changes and demands. Furthermore, environmental communication feedback mechanisms complement short-term memory by providing real-time inputs and interactions within the multi-agent frameworkYogatama et al. (2021). This continuous flow of information allows LLMs to adjust their strategies based on the latest data, facilitating a more agile and effective adaptation process.
By integrating the strengths of both long-term and short-term memory, informed by interdisciplinary research, our LLMs are better positioned to simulate human cognitive processes, offering a more human-like interaction experience. This cross-disciplinary approach not only deepens our understanding of memory system design but also paves the way for future AI research, optimizing LLMs’ performance and adaptability through the emulation of human thought processes.
@@ A. None of the above
A3 (Checker Correction): Add input type checks and raise a ValueError for non-string inputs. Use a specific model for accuracy: distilbert-base-uncased-distilled-squad’. Q4: How to test the play_memory_game functions correctness?
A4 (Assistant): Provides a test function without exception handling.
A4 (Checker Correction): Modify the test function to catch exceptions and provide detailed output for each test case for easier debugging.
Table 25: without prompts
{mdframed}
Q1: How to install the required Python libraries?
A1: Use the following command to install the transformers library:
import subprocess
requirements = [’transformers’]
for package in requirements:
subprocess.run([’pip’, install’, ’-U’, package])
A2: Import the pipeline function to create a question-answering task pipeline:
from transformers import pipeline
A3: Define the play_memory_game function, which takes a context, a question, and the users answer, then checks the correctness using the transformers question-answering pipeline:
def play_memory_game(context, question, user_answer):
raise ValueError(’All arguments must be of type str’)
result = question_answerer(question=question, context=context)
predicted_answer = result[’answer’]
return user_answer.lower() == predicted_answer.lower()
Q4: How to test the play_memory_game function?
A4: The test_play_memory_game function includes three test cases: normal execution, handling non-string input, and incorrect user answer:
def test_play_memory_game():
# Normal execution test
assert play_memory_game(’The capital of France is Paris.’, What is the capital of France?’, Paris’), "Incorrect answer."
# Non-string input test
try:
play_memory_game(123, What is the answer?’, test’)
except ValueError:
pass # Expected failure for non-string input
# Incorrect answer test
assert not play_memory_game(’The Earth revolves around the Sun.’, What does the Moon revolve around?’, Sun’), "Incorrect answer should fail."